TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-05-16 22:04:00
subject: 5\05 Pt 1 ISS On-Orbit Status 05-05-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

06 May 2003

ISS On-Orbit Status 5 May 2003

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.  Week 1 of Increment 7, and Day 10 in space for
Expedition 7 (since 5S launch on 4/26/03).

The crewmembers are back on nominal sleep cycle (2:00am5:30pm EDT)
and have their second off-duty day to help them get "acclimated" to
their new home for the next 185 days.

FE/SO Edward Lu disconnected the UOP (utility outlet panel) power
bypass cable on the RWS DCP (robotics workstation display and control
panel), used for providing SSRMS (space station remote manipulator
system) video support during last weekend's 5S undocking activities.

CDR Yuri Malenchenko conducted the daily checkup of the BIO-5
"Rasteniya-2/Lada-2" zero-G plant growth experiment.

Malenchenko also performed the daily routine maintenance/inspection
of SOZh life support systems, while Ed Lu completed the regular
status checkup of autonomous Lab payloads (PCG-STES010, SAMS, MAMS).

Both crewmembers joined in the weekly planning conference with the
ground via S-band, in preparation for the coming days' activities.

Yuri and Ed also completed their daily physical exercise program, on
TVIS (treadmill with vibration isolation and stabilization), RED
(resistive exercise device), and, for Yuri, on the Russian VELO
ergometer with load trainer

As a long-term item on the Russian task list, Malenchenko is charged
with taking photographs of the PKZ-1V Kromka experiment tablet
deployed on the plume deflector of the SM's plus-pitch thrusters.
[The pictures are taken with the Kodak 760 digital still camera (DSC)
from the EVA hatch 1 window in the DC-1 docking compartment.]

Last Friday's (5/2) replacement of one of the ASV (air supply valves)
filters of the CDRA (carbon dioxide removal assembly) by Bowersox and
Lu was successful.  Inspection of the CDRA pump QD (quick disconnect)
coupling showed no leakage.  The unit was left running for 24 hrs.
for observation, then turned off.  CDRA appears to be entirely
healthy at this time. 

Bowersox and Lu also completed almost all of the planned SSRMS
familiarization activities on Friday, except for one deferred
grapple/ungrapple sequence.

The onboard File Server (FS) is now healthy again, after restoration
of its files.  [There is a question whether full restoration to
Expedition 6 level is really required, with no Expedition 6 present.]

During the undocking activities, a PCS (portable computer system)
laptop failed in the Service Module (SM) and was replaced with an
Airlock (AL) PCS. 

TsUP/Moscow reported that the landing undershoot of the returning
5S/Sojus descent module by about 300 miles on 5/3 was due to an
entirely ballistic reentry of the capsule, performed in (secondary)
Ballistic Descent Mode (BS), instead of the lift-vector-controlled
reentry using banking maneuvers (roll angle changes) commanded by
the (primary) Automatically Controlled Descent Mode (AUS).  An
investigation, with NASA participation, has been initiated.  [Instead
of flying on the closed-loop-guided trajectory designed to reduce
peak deceleration & heating while extending the downrange, the
ballistic mode resulted in a steeper trajectory, ~2g's higher
deceleration forces on the crew (7g max instead of 5g), and an
undershoot of ~300 mi. The crew, now at Star City, was never in any
increased danger, but it took the recovery forces about 4.5 hrs. to
locate and reach them at the backup landing site.]

Today's CEO (crew earth observations) targets, including cities
during daylight and at night (again available as targets in the
current LVLH attitude) were Cairo, Egypt (nadir pass; ESC [electronic
still camera]), Okavango swamp, Botswana (Dynamic event.  As the
present El Nino phase winds down, crew was to try for this sun glint
opportunity to document extent of flooding in the largest swampland
in southern Africa), Miami, Florida (nadir pass; ESC), Navassa
Island, Caribbean (shooting corals around this Caribbean island,
looking a touch right of track), Caracas, Venezuela (nadir pass;
ESC), Brasilia, Brazil (nadir pass; ESC), Mexico City, Mexico (nadir
pass; ESC), Puebla, Mexico (nadir pass; ESC), High Central Andean
Glaciers (shooting any snow or ice pack than can be seen on volcano
summits near nadir), Parana River floods, Argentina (detailed nadir
views of this largest flood in Argentina's history were requested.
Then, 30-60 secs later, crew was to try for high-contrast views
within the sun glint disc [looking directly back along the line of
track]), Puebla, Mexico (nadir pass), El Paso-Las Cruces region
(nadir pass), Osaka, Japan (left of track), Tokyo, Japan (nadir
pass),  and Bombay, India (right of track). http://eol.jsc.nasa.gov
and http://earthobservatory.nasa.gov 

U.S. and Russian Segment Status (as of 1:21pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS):

Elektron O2 generator is powered On (32 Amp mode).  Vozdukh CO2
scrubber is On (manual mode 5).  U.S. CDRA CO2 scrubber is Off.
TCCS (trace contaminant control subsystem) is operating.  MCA (major
constituents analyzer) is operating.  BMP Harmful Impurities unit:
absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air
conditioner SKV-1 is On; SKV-2 is Off. 

SM Working Compartment:  Pressure (mmHg) -- 742; temperature (deg C)
-- 27.2; ppO2 (mmHg) -- data invalid; ppCO2 (mmHg) -- data invalid.
SM Transfer Compartment:  Pressure (mmHg) -- 751; temperature (deg C)
-- 21.4.
FGB Cabin:  Pressure (mmHg) -- 744; temperature (deg C) -- 21.0.
Node:  Pressure (mmHg) -- 739.44; temperature (deg C) -- 22.4
(shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab:  Pressure (mmHg) -- 741.38; temperature (deg C) -- 22.5;
ppO2 (mmHg) -- 167.6; ppCO2 (mmHg) -- 2.4.
Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 741.28; temperature
(deg C) -- 22.4; shell heater temp (deg C) -- 23.3, ppO2 (mmHg) --
n/a; ppCO2 (mmHg) -- n/a.
PMA-1:  Shell heater temp (deg C) -- 23.4
PMA-2:  Shell heater temp (deg C) -- 21.6

(n/a = data not available)

Propulsion System (PS):

Total propellant load available: 3607 kg (7952 lb) as of 5/1
[SM(774) + FGB(2581) + Progress(252)].  (Capability: SM -- 860 kg;
FGB -- 6120 kg). 

Electrical Power Systems (EPS):

Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B
and 4B both in directed position (Blind mode, non solar-tracking).


SM batteries:  Battery #4 is disconnected, in stand-alone ops; all
other batteries (7) are in "Partial Charge" mode.
FGB batteries:  Battery #5 is disconnected; all other batteries (5)
are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.